how would you solve a system of equations like this one
Y=4x+4
Y=-1.5x+6

Answer by stanbon(75887)

solve a system of equations like this one
Y=4x+4
Y=-1.5x+6
-----
Solve:
4x+4 = -1.5x + 6
2.5x = 2
x = 2/(5/2)
x = 4/5
===================
Solve for "y":
y = 4x + 4
y = 4(4/5)+4
y = 16/5 + 20/5
y = 36/5
------------------
cheers,
Stan H.
===================